OpenStudy (anonymous):

Which of the following best describes the solution to the equation below? 6x + 16 = -22 + 6x A. exactly one real solution, B. no real solutions C. infinite real solutions D. exactly one real solution,

OpenStudy (amoodarya):

6x cancel from both sides and 16=-22 so ? no real solutions

OpenStudy (sleepyjess):

First you need to solve for x. 6x + 16 = -22 + 6x 6x - 6x + 16 = -22 + 6x - 6x 16 = -22 Which isn't possible, so there are no real solutions